OpenStudy (anonymous):

Write the first five terms of the arithmetic sequence with the first term a1 and common difference d. a1 = 1 over 13 , d = 3 over 13 A. 1 over 13, -2 over 13, -5 over 13, -8 over 13, -11 over 13 B. 1 over 13, 40 over 13, 79 over 13, 118 over 13, 157 over 13 C. 1 over 13, 4 over 13, 7 over 13, 10 over 13, 13, over 13 D. 1 over 14, 1 over 16, 1 over 19, 1 over 22, 1 over 25

OpenStudy (kirbykirby):

C: If the common difference is positive, the sequence will grow towards infinity (it will keep growing and growing). We see that 1/13 + 3/13 = 4/13 Then 4/13 + 3/13 = 7/13 Then 7/13 + 3/13 = 10/13 Then 10/13 + 3/13 = 13/13 This is in contrast to answer A, which is similar, except the trend is going the other way (1/13 - 3/13 = -2/13, then -2/13 - 3/13 = -5/13,...) but the common difference is positive so you choose the sequence growing towards larger numbers.

OpenStudy (anonymous):

so the answer is C or A ?

OpenStudy (kirbykirby):

C, because the sequence is getting larger and larger. If you are still asking qhat the answer is after this, I don't think you still understand the question?

OpenStudy (kirbykirby):

what*